if cos3x/cosx = 1/3 (0<=x<=90)
what is the value of sin3x/sinx ?

Grade:11

2 Answers

shubhra pathak
13 years ago

cos3x/cosx= (4cos^3 x-3cosx)/cosx=1/3

=> 4cos^2 x-3=1/3

cos^2 x=5/6

as we know that cos^2 x+sin^2 x=1

so,  sin^2 x=1/6

and    sin3x/sinx= (3sinx-4sin^3 x)/sinx

=> 3-4sin^2 x=7/3

asnwer>> 7/3

Shreya Chandwadkar
13 years ago

thank you !! but i could solve it.... by the same method after posting it!!